OpenStudy (anonymous):

What is the sum of a 7-term geometric series if the first term is -11, the last term is -45,056, and the common ratio is -4?

OpenStudy (anonymous):

I'm sorry, but I have to ask again: Do you know the formula?

OpenStudy (anonymous):

is it the same formula lol..

OpenStudy (anonymous):

not like arithmetic ... O_O

OpenStudy (anonymous):

this seems interesting... let me try it on paper, I'll let you know IF i can do it ;)

OpenStudy (anonymous):

haha kk thanx

OpenStudy (anonymous):

uff... this is not easy (for me) ... using logs It seems it doesn't work. Maybe I made a mistake somewhere... Let me think :(

OpenStudy (anonymous):

... I can't, is just impossible. Do you have any advanced calculator? to find "n" here.. \[\LARGE \log_4(n)+1-n=\log_4(11)\]

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

.... well I guess I can't do it. I have to confess that this is a cute problem , but no matter how much I beg her, she just won't give up :) lol

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

call someone maybe they can help :S ... and I think that one with logs there above, I just broke some rules lol. now I see it, my math sucks anyway.

OpenStudy (anonymous):

Shouldn't we use the formula:\[a \frac{1-r^n}{1-r}\]

OpenStudy (anonymous):

yes - you dont need to know the last term

OpenStudy (anonymous):

if you don't know the last term what n is going to be ? :O

OpenStudy (anonymous):

the formula is: \[\LARGE Sn=a_1\cdot {1-r^n\over 1-r }\]

OpenStudy (anonymous):

S7 = -11 * (1 - (-4)^7) --------- 1-(-4)

OpenStudy (anonymous):

Geometric series shall be in the form of a, ar, ar^2 , ar^3,....ar^6 (these are first 7 terms) a=-11 ar^6=-45056 taking the ratio ar^6/a=-45056/-11 r^6=4096 (as given in the question common ratio r=4 will satisfy this). Just to make sure everything is in geometric sum of first 7 terms gives (I am not using striaght formula, I shall derive one that helps you to verify) y=a(1+r+r^2+r^3+...+r^6) y/a=1+r+r^2+r^3+...+r^6 --- (1) now mulitplying with r on both sides we get ry/a=r +r^2+r^3+...+r^7 --- (2) subtracting (1) from (2) ry/a-y/a=r^7-1 (remaining terms cancel each other) y(r-1)/a=r^7-1 y=a(r^7-1)/(r-1) [r should not be equal to 1] in general sum of n terms in geometric series is a(r^(n+1)-1)/(r-1) [considering r>1] a(1-r^(n+1))/(1-r) [considering r<1] sum of 7 terms with a=-11, r=4 gives -11*(4^7-1)/(4-1) -11*(16384-1)/3 -11*(16383/3) -11*5461 -60071

OpenStudy (anonymous):

The common ratio is -4, not 4.

OpenStudy (anonymous):

Mine yielded -36047, but I am sure I made a mistake.

OpenStudy (anonymous):

ok theses are the 2 questions and the possible answer choices if it helps you guys--->What is the sum of a 7-term geometric series if the first term is -11, the last term is -45,056, and the common ratio is -4? (3 points) A) -143,231 B) -36,047 C) 144,177 D) 716,144 21. What is the sum of an 8-term geometric series if the first term is -11, the last term is 180,224, and the common ratio is -4? (3 points) A) -143,231 B) -36,047 C) 144,177 D) 716,144

OpenStudy (anonymous):

AAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAWWWWWWWWWWWWWWWWWWWWWWWWWW :@:@:@ I SWEAR I GONN' BREAK MY PC RIGHT NOW :@ AAAHHH... (angry ) how come I didn't know it :(

OpenStudy (anonymous):

I got B from the first.

OpenStudy (anonymous):

I didn't read the question properly, and I thought that we have to find a_n then we have to find S_n and that "n" I just couldn't find . :(

OpenStudy (anonymous):

i just want to say thank you guys so much for all the help and im srry if its making u mad :/

OpenStudy (anonymous):

[-11 * 16385] / 5 = - 36047

OpenStudy (anonymous):

And I got C for the second.

OpenStudy (anonymous):

thanks bmp :)

OpenStudy (anonymous):

\[\Large S_8=-11\cdot {1-(-4)^8 \over 1-(-4)} \] pff.. Peace of cake .. Lord I'm stupid :(

OpenStudy (anonymous):

No problem, mate :-)

OpenStudy (anonymous):

Yeah, its my bad. common ratio is -4 instead of 4 :) That makes the same as bmp answer :)